# Technology, Institution, and Regional Growth: Evidence from Mineral Mining Industry in Industrializing Japan

### v14

@@ Abstract
Coal extraction was an influential economic activity in interwar Japan. Initially, coal mines employed both males and females as the workforce in the pits. However, the innovation of labor-saving technologies and the renewal of traditional extraction methodology induced institutional change. This was manifested by the revision of labor regulations affecting female miners in the early 1930s. This dramatically changed the mining workplace, making skilled males the principal miners engaged in underground work. This paper investigates the impact of coal mining on regional growth and assesses how the institutional changes induced by the amended labor regulations affected its processes. By linking the mines’ location information with both registration and census-based statistics, it was found that coal mines led to remarkable population growth. The labor regulations accelerated local population growth by forcing female miners to exit the labor market and form families. The regulations prohibited female miners from risky underground work. This reduction in occupational hazard also improved early-life mortality via the mortality selection mechanism in utero.
@@ 1 Introduction
This study is relevant to the broader literature on the economic impact of resource extraction. Predictions regarding agglomeration theory imply that resource abundance can increase the demand for local labor in the mining sector and related industries, thereby causing an influx of population and agglomeration (Rosenthal and Strange 2004; Duranton and Puga 2004; Lederman and Maloney 2007). [^2] Induced agglomeration can improve efficiency in the local labor markets through, for instance, better matching and learning. This can have spillover effects in the non-mining sectors (Black et al. 2005; Bjørnland and Thorsrud 2016). See Corden and Neary (1982); Eastwood and Venables (1982); Corden (1984); van Wijnbergen (1984); Neary and van Wijnbergen (1984) for earlier theoretical studies on the impact of resource discoveries. However, growth of the resource extraction industry may stagnate the development of non-resource extraction industries by reallocating resources from the non-resource tradable sector (Sachs and Warner 1999, 2001). [^3] The wealth effect mechanism suggests that resource abundance increases the demand for commodities, thereby reallocating resources from the tradable to non-tradable sector and increasing the imports of tradables. This can reduce labor supply, leading to greater earnings in the local economy. The direct effect on the relative demand of extraction firms and workers is higher prices of personal and business services. Other political-economic factors that may disrupt the growth of the non-resource sector are rent-seeking and corruption. Caselli and Guy (2009, 2013) summarize these theoretical implications. See Papyrakis and Gerlagh (2007) for empirical evidence using United States data. There is still debate on whether proximity to coal mines was necessary for regional development during Europe’s industrialization. A recent study by Fernihough and O’Rourke (2020) found that after the mid-eighteenth century, European cities closer to coalfields had larger populations than other cities. [^4] Similarly, some economic history studies have argued for the importance of better access to coalfields in subsequent industrial development (Mathias 2001; Pollard 1981). By contrast, Mokyr (1977) observed that the distribution of coal supplies did not contribute to industrialization. My main results support the former assertion by showing that the coal mines led to population growth in the local economy. Importantly, this study contributes to the literature analyzing the mechanism behind population growth due to resource extraction. [^5] Although studies have analyzed the impact of resource abundance on population growth, the mechanisms underlying this relationship have been understudied. See for example Black et al. (2005); Michaels (2010); Fernihough and O’Rourke (2020). I demonstrate that while the population growth in the mining area initially occurred due to internal migration, the miner households formed families in response to the institutional change.
@@ 2.1 Coal Mining Industry
The Mining Code of 1892 allowed private companies to have mining concessions, accelerating the installation of modern technologies in mines and the development of Japan’s mining sector. Winding engines were installed in coal and copper mines in the 1890s, and the flame-smelting process was adopted in gold and silver mines in the 1900s (Ishii 1991, pp. 220–221). Before the First World War (WWI), however, the Japanese coal industry had developed as a labor-intensive export industry. Then, the outbreak of WWI increased miners’ average wage rates, reducing the coal industry’s comparative advantage. Figure 1 illustrates the development of the coal industry and shows the negative shock in the early 1920s. This motivated coal mining firms to reduce the number of miners and adopt new extraction machines to improve productivity. [^9] See Sumiya (1968, pp. 387–393) for details on the machinery adoption process. Mining firms reallocated resources to relatively productive mines to improve their average productivity. Consequently, average labor productivity in the coal mining sector increased from the 1920s, particularly in the 1930s (Okazaki 2021). [^10] Generally, during the interwar period, the mining and manufacturing industries experienced steady capital investment, particularly in production facilities and machines, and public investments (Nakamura 1971, pp. 138–143). Consequently, the relative contribution of manufacturing and mining, which provided energy and resources to manufacturing, was the largest at 42-51% among all economic sectors during this period (Minami 1994, p.90). Figure 1 shows that total coal production increased during the interwar period. [^11] As the moderate decline in the early 1930s confirmed, the impact of the depression was not substantial compared to the other countries. See Online Appendix A.1 for details.
@@
2.2 Mining Technology, Occupational Hazards, and Regulations of Female Miners
@@
By the end of the 19th century, a large proportion of mines had completed mechanizing the hauling processes. This was done by installing equipment such as drainage pumps and hoisting machines. However, this did not mean that the miners’ workload related to hauling operations was reduced. Instead, such mechanization in the main shafts increased the workload of miners extracting and transporting coal around the coalface (Sumiya 1968, pp. 310–312). Moreover, the environment in the pit was generally poor. The air was polluted by dust, oxygen levels were low, and there was a risk of carbon monoxide poisoning. There was also a high fire risk from dynamite blasting and spontaneous combustion of gas and coal. Accidents caused by rockfall occurred frequently. Available statistics indicate that the average number of accidents per year in the pits between $1917$ and $1926$, causing miners’ injuries and deaths, was $142,595$, of which $61,112$ were from rockfall. Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by rockfall, but they still occurred $311$ times per year on average between $1917$ and $1926$. [^13] These statistics are from Japan Mining Association (1928, p. table.6).
Notably, while miners included single male workers, many of them were married couples. This was because miners usually worked in teams of two to three (Sumiya 1968, pp. 300–308). The male skilled miner, called saki yama, extracted coal at the face; meanwhile, the female miner, named ato yama, carried that coal to the coal wagon at the gangway. She traveled through steep pits carrying bamboo baskets called sura (Sumiya 1968, p. 319). As coal is a bulky resource, the heavy work burden on pregnant women could have increased the risk of adverse pregnancy outcomes (Ahmed and Jaakkola 2007). Miners worked in pits with dirty air containing particulates, which eventually damaged their lungs and increased the risk of respiratory diseases such as tuberculosis (Sumiya 1968, pp.  410–411). [^14] A miners’ health survey found that coal miners had the highest respiratory disease prevalence among all types of miners at that time, which indicated how hard the work in the coal mines was (Japan Mining Association 1928, p. 3). Relatedly, a growing body of medical literature has revealed that particulates and radionuclides from coal increase the risk of stillbirths and premature deaths due to mothers’ respiratory diseases (Landrigan et al. 2017; Lin et al. 2013).
@@
An important event that changed the harsh labor pattern experienced by female miners was an institutional change occurring in the early 1930s. Following the International Labor Conference after WWI, the Social Affairs Bureau held a consultation with mining companies in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Initially, mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night. However, the renewal of mining methods driven by new extraction technologies changed the attitudes of the mining companies in the late 1920s. The transition from the traditional room-and-pillar mining method to the longwall method was accompanied by the adoption of technologies such as coal cutters and conveyors. Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of those machines (Nishinarita 1985, p. 91). Specifically, the conveyors brought extracted coals from pits to the gangway more efficiently than the female miners, strongly incentivizing the mining companies to prohibit underground work by female miners (Tanaka and Ogino 1977, p. 79). [^15] Online Appendix A.2 provides finer details in this mechanization process. Online Appendix A.3 provides cross-sectional evidence on the replacement of female miners by machines. Finally, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work by female miners. As the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930. This change induced by mechanization was commonly observed in Japanese coal mines (Tanaka 1984, p. 392). [^16] A detailed chronology of technology adoptions in each coal mine in Chikuhō coalfield confirms this movement (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973).
@@
2.3 Mining Technology and Pollutions
@@
First was the soot and smoke caused by coal burning to run steam-powered winding machines. Steam-powered winding engines were introduced early in the mechanization of the hauling process (Section 2.1) and exhausted smoke generated from burning coal in the boiler. However, these steam-powered winders were gradually replaced by electric winders during the interwar period (Japan Mining Association 1930, 1931). Thus, this channel may be unlikely to impact the regional population health. Second, presumably more relevant, is pollution from wastewater generated during mining. As coal mining became increasingly mechanized, some mines began introducing water-washing machines during coal sorting. Unfortunately, there is little historical documentation systematically describing pollution caused by wastewater from coal mines. However, coal sludge from coal selection started contaminating rivers at the time. An example is Onga river in the Chikuhō coalfield, which was called a “zenzai” (sweet red-bean soup) because the water was contaminated by the liquid waste (Chiba and Yamada 1964). Coal sludge contains several metal toxins, including mercury (Hg) and cadmium (Cd). These are known to be associated with the incidence of miscarriages and stillbirths (Amadi et al. 2017; Sergeant et al. 2022). This suggests the possibility that early-life mortality would have been higher in the coal mining regions with greater river accessibility. From the 1920s, however, some mines installed settling ponds to purify wastewater and recirculate it during coal cleaning (Advisory Council on Mining 1932, pp. 313–316). This could mitigate the pollution risks around the mines due to the wastewater.
@@ 3 Conceptual Framework
The regulation of female miners’ working conditions described in the previous section would influence the fertility in the mining area. The decline in females’ relative wage rate due to regulation can decrease the opportunity cost of housework, whereas it can also decrease their earnings in the labor market. While the former mechanism could lead to higher fertility through the substitution effect from consumption, the latter may result in lower fertility via the income effect. This section offers a brief overview of a theoretical framework predicting the relevant mechanism given the historical evidence about the coal mines in prewar Japan. To do so, I consider the collective model proposed by Siegel (2017) that illustrates the optimal fertility based on the couple household’s utility maximization. [^18] For the sake of brevity, a simplified model that abstains the utility from leisure and home production parameters is considered. This offers a materially similar interpretation of the optimal fertility choice to the original generalized model. See Browning et al. (2014) for a comprehensive review on the cooperative and non-cooperative models of the couples.
@@
Online Appendix B provides details of the derivation. Under a paternalistic social norm stipulated in the Old Civil Code, the head’s housework productivity, particularly childrearing, was negligible ($\kappa_{m}=0$). It follows that condition 1 implies $\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}$ for my empirical setting. Since the marketization of childcare ($\gamma$) was considerably low at that time, the optimal fertility was also negatively associated with the relative wage ($w_{f}/w_{m}$). [^19] As mechanization progressed, miners were required to accumulate knowledge and experience in machinery operation. Management, therefore, had an incentive to encourage miners to form families and continue employment over a long term (Ogino 1993). Thus, many coal mines were equipped with nurseries and other welfare facilities from the 1920s (Ministry of Commerce and Industry 1926). However, evidence shows that these nurseries required fees and had covered only a small part of their childcare requirements. See Online Appendix B for details. This result suggests the following two predictions:
@@ 4.1 Mine Deposit
I use official reports named Zenkoku kōjyō kōzan meibo (lists of factories and mines) published by Association for Harmonious Cooperation (1932, 1937) (hereafter, called the AHC), which document coal mine locations measured in October 1931 and October 1936. The AHC listed all mines employing $50$ miners and more. Thus, the main target of this study is the average impact of small- to large-scale coal mines and not the micro-scale collieries employing fewer than $50$ mimers. [^22] Generally, several coal mines are clustered in a coal mining municipality. For example, 85 (108) out of 93 (115) municipalities with mines had one to four mines in the 1930 (1935) sample. In this sense, this study estimates the average impact of these general coal mining municipalities. In Online Appendix D.11, I confirmed that the main findings are not sensitive to a small number of municipalities with several mines. Despite this, the AHC is still comprehensive as it includes information on small-scale mines, which have been neglected in the literature.
@@
Figure 2 indicates the coal mine locations from the AHC and the corresponding treatment and control groups by year. Most coal mines are located in three specific areas in the southernmost, central, and northeastern regions. These are representative coalfields called Chikuhō, Jyōban, and Ishikari-Kushiro Coalfields, respectively. Notably, a comparison of Figures 2a and 2b shows that these agglomerations were stable over time. For a closer look, Figure 3 provides an example map for the most representative coalfield called Chikuhō Coalfield in Kyūshū, Japan’s southernmost island (indicated in Figure 2a). One can observe that the distribution of mines has changed little, with only a few mines either opened or closed between 1931 and 1936. This feature of coal mines makes it difficult to use within-variation in the spatiotemporal distribution of the mines for identification. Online Appendix D.8 discusses this point in detail. [^25] In short, it is not feasible to implement a difference-in-differences estimator for the full sample because of the lack of within-variation (i.e., the number of municipalities that experienced the opening/closing of the coal mines). Nevertheless, in Online Appendix D.8, the results from the two-group by two-period difference-in-differences setting for a small subsample including Ogawa and Miyoshiyama coal mines will be demonstrate (Figure 2b), which provide the results reasonable to the findings.
Finally, the generation of the analytical samples will be explained. From $11,151$ total municipalities, the municipalities within a $30$ km radius from the centroid of the municipalities with mines are first retained. The intersection of the different types of mines are then excluded, such as municipalities within $5$ km from a gold, silver, and copper (GSC) mine. Online Appendix Table C.3 presents the summary statistics of the treatment variables. The number of municipalities in the analytical sample based on the 1931 and 1936 AHC were $1,140$ and $1,364$, respectively. Meanwhile, the proportion of treated municipalities was approximately $15$% and $13$%, respectively, for each sample. [^26] The number of municipalities with coal mines in each analytical sample is 93 and 115, respectively. The slight reduction in the share of the treatment group is due to the increase in the number of enclaves, which is explained above. A mine opening in the enclave area increases the number of controlled (i.e., surrounding) municipalities, which leads to a decrease in the relative share of the treatment group. Examples are the Ogawa and Miyoshiyama coal mines in Tōhoku region (Figure 2b).

### v15

@@ Abstract
Coal extraction was an influential economic activity in interwar Japan. Initially, coal mines employed both males and females as the workforce in the pits. However, the innovation of labor-saving technologies and the renewal of traditional extraction methodology induced institutional change. This was manifested by the revision of labor regulations affecting female miners in the early 1930s. This dramatically changed the mining workplace, making skilled males the principal miners engaged in underground work. This paper investigates the impact of coal mining on regional growth and assesses how the institutional changes induced by the amended labor regulations affected its processes. By linking the mines’ location information with both registration and census-based statistics, it was found that coal mines led to remarkable population growth. Fertility rate increased following the implementation of labor regulations that required female miners to leave the workforce and start families. The regulations prohibited female miners from risky underground work. This reduction in occupational hazard also improved early-life mortality via the mortality selection mechanism in utero.
@@ 1 Introduction
This study is relevant to the broader literature on the economic impact of resource extraction. Predictions regarding agglomeration theory imply that resource abundance can increase the demand for local labor in the mining sector and related industries, thereby causing an influx of population and agglomeration (Rosenthal and Strange 2004; Duranton and Puga 2004; Lederman and Maloney 2007). [^2] Induced agglomeration can improve efficiency in the local labor markets through, for instance, better matching and learning. This can have spillover effects in the non-mining sectors (Black et al. 2005; Bjørnland and Thorsrud 2016). See Corden and Neary (1982); Eastwood and Venables (1982); Corden (1984); van Wijnbergen (1984); Neary and van Wijnbergen (1984) for earlier theoretical studies on the impact of resource discoveries. However, growth of the resource extraction industry may stagnate the development of non-resource extraction industries by reallocating resources from the non-resource tradable sector (Sachs and Warner 1999, 2001). [^3] The wealth effect mechanism suggests that resource abundance increases the demand for commodities, which reallocates resources from the tradable to the non-tradable sector and boosts the imports of tradables. This shift can decrease labor supply, leading to higher earnings in the local economy. The direct effect on the relative demand of extraction firms and workers is an increase in the prices of personal and business services. Additionally, other political-economic factors, such as rent-seeking and corruption, may disrupt the growth of the non-resource sector. Caselli and Guy (2009, 2013) summarize these theoretical implications. For empirical evidence using data from the United States, see Papyrakis and Gerlagh (2007). There is still debate on whether proximity to coal mines was necessary for regional development during Europe’s industrialization. A recent study by Fernihough and O’Rourke (2020) found that after the mid-eighteenth century, European cities closer to coalfields had larger populations than other cities. [^4] Similarly, some economic history studies have argued for the importance of better access to coalfields in subsequent industrial development (Mathias 2001; Pollard 1981). By contrast, Mokyr (1977) observed that the distribution of coal supplies did not contribute to industrialization. My main results support the former assertion by showing that the coal mines led to population growth in the local economy. Importantly, this study contributes to the literature analyzing the mechanism behind population growth due to resource extraction. [^5] Although studies have analyzed the impact of resource abundance on population growth, the mechanisms underlying this relationship have been understudied. See for example Black et al. (2005); Michaels (2010); Fernihough and O’Rourke (2020). I demonstrate that while the population growth in the mining area initially occurred due to internal migration, the miner households formed families in response to the institutional change.
@@ 2.1 Coal Mining Industry
The Mining Code of 1892 allowed private companies to have mining concessions, accelerating the installation of modern technologies in mines and the development of Japan’s mining sector. Winding engines were installed in coal and copper mines in the 1890s, and the flame-smelting process was adopted in gold and silver mines in the 1900s (Ishii 1991, pp. 220–221). Before the First World War (WWI), however, the Japanese coal industry had developed as a labor-intensive export industry. Then, the outbreak of WWI increased miners’ average wage rates, reducing the coal industry’s comparative advantage. Figure 1 illustrates the development of the coal industry and shows the negative shock in the early 1920s. This motivated coal mining firms to reduce the number of miners and adopt new extraction machines to improve productivity. [^9] See Sumiya (1968, pp. 387–393) for details on the machinery adoption process. Mining firms reallocated resources to relatively productive mines to improve their average productivity. Consequently, average labor productivity in the coal mining sector increased from the 1920s, particularly in the 1930s (Okazaki 2021). [^10] Generally, during the interwar period, the mining and manufacturing industries experienced steady capital investment, particularly in production facilities and machines, and public investments (Nakamura 1971, pp. 138–143). Consequently, the relative contribution of manufacturing and mining, which provided energy and resources to manufacturing, was the largest at 42-51% among all economic sectors during this period (Minami 1994, p.90). Figure 1 shows that total coal production increased during the interwar period. [^11] As the moderate decline in the early 1930s confirmed, the effect of the depression was not substantial compared to that in other countries. For details, see Online Appendix A.1.
@@
2.2 Mining Technology and Labor Patterns
@@
By the end of the 19th century, most mines had mechanized the hauling processes in the main shaft. This was done by installing equipment such as drainage pumps and hoisting machines. However, this did not mean that the miners’ workload related to hauling operations was reduced. Instead, such mechanization in the main shafts increased the workload of miners extracting and transporting coal around the coalface (Sumiya 1968, pp. 310–312). Notably, while miners included single male workers, many of them were married couples. This was because miners usually worked in teams of two to three (Sumiya 1968, pp. 300–308). The male skilled miner (saki yama) extracted coal at the face; meanwhile, the female miner (ato yama) carried that coal to the coal wagon at the gangway. She traveled through steep pits while carrying wooden baskets (Sumiya 1968, p. 319). In addition, the environment in the pit was generally poor. Accidents caused by rockfall occurred frequently. [^13] Available statistics indicate that the average number of accidents per year in the pits between $1917$ and $1926$, causing miners’ injuries and deaths, was $142,595$, of which $61,112$ were from rockfall. The air was polluted by dust, oxygen levels were low, and there was a risk of carbon monoxide poisoning. There was also a fire risk from dynamite blasting and spontaneous combustion of gas and coal. [^14] Accidents caused by explosions, poisoning, and suffocation were fewer than those caused by rockfall, but they still occurred $311$ times per year on average between $1917$ and $1926$. These accident statistics are from Japan Mining Association (1928, p. table.6).
@@
An institutional change that transformed the harsh working conditions of female miners took place during the interwar period. Following the International Labor Conference after WWI, the Social Affairs Bureau held a consultation with mining companies in the 1920s to revise the Miners’ Labor Assistance Regulations of 1916. Initially, mining companies, who relied heavily on female miners for production, opposed the Social Affairs Bureau’s insistence on prohibiting women from working underground and late-night. However, the renewal of the mining method changed the attitudes of mining companies in the late 1920s (Online Appendix A). In fact, the transition from the traditional room-and-pillar mining method to the longwall method was accompanied by the adoption of technologies such as coal cutters and conveyors. Compared to the room-and-pillar method, the coal extraction space of the longwall method is larger, allowing for the introduction of those machines (Nishinarita 1985, p. 91). Specifically, the conveyors brought extracted coals from pits to the gangway more efficiently than the female miners, strongly incentivizing the mining companies to prohibit underground work by female miners (Tanaka and Ogino 1977, p. 79). [^15] Online Appendix A.2 provides finer details in this mechanization process. Online Appendix A.3 provides cross-sectional evidence on the replacement of female miners by machines. Finally, the Revised Miners’ Labor Assistance Regulations of September 1928 prohibited some underground and late-night work by female miners. As the grace period for enforcement of the regulations lasted until 1933, the number of female miners gradually declined from around 1930. This change induced by mechanization was commonly observed in Japanese coal mines (Tanaka 1984, p. 392). [^16] A detailed chronology of technology adoptions in each coal mine in Chikuhō coalfield confirms this movement (Compilation Committee of the Chronology of Chikuhō Coal Mining Industry 1973).
@@
Although systematic statistics are unavailable, unemployed female miners were viewed as likely to become housewives. Iwaya (1997, p. 11) describes this shift, stating, “the mining community gave women a new role, not in labor, but in supporting labor in the home.” The official report of the 1930 census shows that the female labor force participation rate in the mining area was significantly lower than in the surrounding agricultural municipalities (Online Appendix D.5). Although some female workers could find employment in the domestic sector, the average wage rate was lower than in the mining sector. Consequently, this movement worsened the relative wage rate in the mining area. Additionally, the official report of the population census indicates that the age distribution among female miners did not change following the revised regulations (Online Appendix A.4). [^17] The official census report showed that most (more than 70%) of female miners were married, even after the establishment of the revised regulation (Statistics Bureau of the Cabinet 1935b, pp. 160–161). This figure is similar to that in 1920, meaning that the proportion of marriages among female miners was stable during this period. This suggests that the overall productivity in the mining area was stable, before and after the revised regulation.
@@
2.3 Pollution and Occupational Hazards for Female Miners
Under the traditional coal extraction system, female miners carried the extracted coal to the wagon. As coal is a bulky resource, the heavy work burden on pregnant women could have increased the risk of adverse pregnancy outcomes (Ahmed and Jaakkola 2007). In addition, miners worked in pits with dirty air containing particulates, which eventually damaged their lungs and increased the risk of respiratory diseases such as tuberculosis (Sumiya 1968, pp. 410–411). A miners’ health survey found that coal miners had the highest respiratory disease prevalence among all types of miners at that time, which indicated how hard the work in the coal mines was (Japan Mining Association 1928, p. 3). A growing body of medical literature has revealed that particulates and radionuclides from coal increase the risk of stillbirths and premature deaths due to mothers’ respiratory diseases (Landrigan et al. 2017; Lin et al. 2013).
@@ 2.3 Pollution and Occupational Hazards for Female Miners
First was the soot and smoke caused by coal burning to run steam-powered winding machines. Steam-powered winding engines were introduced early in the mechanization of the hauling process (Section 2.1) and exhausted smoke generated from burning coal in the boiler. However, these steam-powered winders were gradually replaced by electric winders during the interwar period (Japan Mining Association 1930, 1931). Second, presumably more relevant, is pollution from wastewater generated during mining. As coal mining became increasingly mechanized, some mines began introducing water-washing machines during coal sorting. Unfortunately, there is little historical documentation systematically describing pollution caused by wastewater from coal mines. However, coal sludge from coal selection started contaminating rivers at the time. An example is Onga river in the Chikuhō coalfield, which was called a “zenzai” (sweet red-bean soup) because the water was contaminated by the liquid waste (Chiba and Yamada 1964). Coal sludge contains several metal toxins, including mercury (Hg) and cadmium (Cd). These are known to be associated with the incidence of miscarriages and stillbirths (Amadi et al. 2017; Sergeant et al. 2022). This suggests the possibility that early-life mortality would have been higher in the coal mining regions with greater river accessibility. From the 1920s, however, some mines installed settling ponds to purify wastewater and recirculate it during coal cleaning (Advisory Council on Mining 1932, pp. 313–316). This could mitigate the pollution risks around the mines due to the wastewater.
@@ 3 Conceptual Framework
The regulation of female miners’ working conditions described in the previous section would influence the fertility in the mining area. Regulations that lead to a decrease in women’s relative wage rates can reduce the opportunity cost of housework, while also lowering their earnings in the labor market. While the former mechanism could lead to higher fertility through the substitution effect from consumption, the latter may result in lower fertility via the income effect. This section offers a brief overview of a theoretical framework predicting the relevant mechanism given the historical evidence about the coal mines in prewar Japan. To do so, I consider the collective model proposed by Siegel (2017) that illustrates the optimal fertility based on the couple household’s utility maximization. [^19] For the sake of brevity, a simplified model that abstains the utility from leisure and home production parameters is considered. This offers a materially similar interpretation of the optimal fertility choice to the original generalized model. See Browning et al. (2014) for a comprehensive review on the cooperative and non-cooperative models of the couples.
@@
Online Appendix B provides details of the derivation. Under a paternalistic social norm stipulated in the Old Civil Code, the head’s housework productivity, particularly childrearing, was negligible ($\kappa_{m}=0$). [^20] The same predictions can be derived under a more plausible assumption, $\kappa_{m}\simeq 0$ (Online Appendix B). It follows that condition 1 implies $\frac{w_{f}}{w_{m}}<\frac{1-\gamma}{\gamma}$ for my empirical setting. Since the marketization of childcare ($\gamma$) was considerably low at that time, the optimal fertility was also negatively associated with the relative wage ($w_{f}/w_{m}$). [^21] As mechanization progressed, miners were required to accumulate knowledge and experience in machinery operation. Management, therefore, had an incentive to encourage miners to form families and continue employment over a long term (Ogino 1993). Thus, many coal mines were equipped with nurseries and other welfare facilities from the 1920s (Ministry of Commerce and Industry 1926). However, evidence shows that these nurseries required fees and had covered only a small part of their childcare requirements. See Online Appendix B for details. This result suggests the following two predictions:
@@ 4.1 Mine Deposit
I use official reports named Zenkoku kōjyō kōzan meibo (lists of factories and mines) published by Cooperative Society (1932, 1937) (hereafter, called the CS), which document coal mine locations measured in October 1931 and October 1936. The CS listed all mines employing $50$ miners and more. Thus, the main target of this study is the average impact of small- to large-scale coal mines and not the micro-scale collieries employing fewer than $50$ mimers. [^24] Generally, several coal mines are clustered in a coal mining municipality. For example, 85 (108) out of 93 (115) municipalities with mines had one to four mines in the 1930 (1935) sample. In this sense, this study estimates the average impact of these general coal mining municipalities. In Online Appendix D.12, I confirmed that the main findings are not sensitive to a small number of municipalities with several mines. Despite this, the CS is still comprehensive as it includes information on small-scale mines, which have been neglected in the literature.
@@
Figure 2 indicates the coal mine locations from the CS and the corresponding treatment and control groups by year. Most coal mines are located in three specific areas in the southernmost, central, and northeastern regions. These are representative coalfields called Chikuhō, Jyōban, and Ishikari-Kushiro Coalfields, respectively. Notably, a comparison of Figures 2a and 2b shows that these agglomerations were stable over time. For a closer look, Figure 3 provides an example map for the most representative coalfield called Chikuhō Coalfield in Kyūshū, Japan’s southernmost island (indicated in Figure 2a). One can observe that the distribution of mines has changed little, with only a few mines either opened or closed between 1931 and 1936. This feature of coal mines makes it difficult to use within-variation in the spatiotemporal distribution of the mines for identification. Online Appendix D.9 discusses this point in detail. [^27] In short, it is not feasible to implement a difference-in-differences estimator for the full sample because of the lack of within-variation (i.e., the number of municipalities that experienced the opening/closing of the coal mines). Nevertheless, in Online Appendix D.9, the results from the two-group by two-period difference-in-differences setting for a small subsample including Ogawa and Miyoshiyama coal mines will be demonstrate (Figure 2b), which provide the results reasonable to the findings.
Finally, the generation of the analytical samples will be explained. From $11,151$ total municipalities, the municipalities within a $30$ km radius from the centroid of the municipalities with mines are first retained. The intersection of the different types of mines are then excluded, such as municipalities within $5$ km from a gold, silver, and copper (GSC) mine. Online Appendix Table C.3 presents the summary statistics of the treatment variables. The number of municipalities in the analytical sample based on the 1931 and 1936 CS were $1,140$ and $1,364$, respectively. Meanwhile, the proportion of treated municipalities was approximately $15$% and $13$%, respectively, for each sample. [^28] The number of municipalities with coal mines in each analytical sample is 93 and 115, respectively. The slight reduction in the share of the treatment group is due to the increase in the number of enclaves, which is explained above. A mine opening in the enclave area increases the number of controlled (i.e., surrounding) municipalities, which leads to a decrease in the relative share of the treatment group. Examples are the Ogawa and Miyoshiyama coal mines in Tōhoku region (Figure 2b).